Suppose that the walking step lengths of adult males are normally distributed with a mean of 2.4 feet and a standard deviation of 0.3 feet. A sample of 34 men’s step lengths is taken. (a) Find the probability that an individual man’s step length is less than 2.1 feet. (b) Find the probability that the mean of the sample taken is less than 2.1 feet.
Answer #1

(a) The probability that an individual man’s step length is less than 2.1 feet = 0.1587

(b) The probability that the mean of the sample taken is less than 2.1 feet = 0.0000

M = 2.4 , 0 = 0.3, n = 34 0,3 @ PCX < 2-1) = P( X =4< 2.1-2.4) = P(Z <-1) = [0.1587 @ PLX < 2-1) = PO < 20.05.2014) = P(Z <-5